OpenStudy (anonymous):

Let \( \pi(n)\) be the number of primes less or equal to n. Show that \[ n^{\pi(2n)-\pi(n)}<4^{n} \]

OpenStudy (anonymous):

sir i tried induction..but no result are we supposed to prove by induction or not?

OpenStudy (anonymous):

The proof I know does not need induction.

OpenStudy (anonymous):

@mukushla got any clue?

OpenStudy (anonymous):

nope :(

OpenStudy (anonymous):

Hint \[ 4^n=(1+1)^{2n}> {2n \choose n} \]

OpenStudy (anonymous):

for \(n\le p_k\le 2n\)\[\prod p_k | \frac{(n+1)(n+2)...(2n)}{n!}={2n \choose n} \Rightarrow\prod p_k <{2n \choose n}\]\[n^{\pi(2n)-\pi(n)}< \prod p_k<\binom{2n}{n}< 2^{2n}=4^{n}\]